OpenStudy (anonymous):

Which of the following is the slope between the two points (1, 5) and (6, 4)? a.1/5 b.- 1/5 c. 5 d.-5

OpenStudy (anonymous):

rise over run = (5-4)/? Can you find the denominator?

OpenStudy (anonymous):

ummm no im not good in slope

OpenStudy (saifoo.khan):

\[Slope = \frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (saifoo.khan):

\[Slope = \frac{4-5}{6-1}\]Solve,

OpenStudy (anonymous):

it it c

OpenStudy (saifoo.khan):

Nope.

OpenStudy (anonymous):

it b

OpenStudy (saifoo.khan):

How? :O

OpenStudy (anonymous):

im not sure

OpenStudy (anonymous):

can u please explian it to me

OpenStudy (anonymous):

Babyzoe, did you see the formula above?

OpenStudy (saifoo.khan):

http://www.twiddla.com/709951

OpenStudy (anonymous):

yea i saw it

OpenStudy (saifoo.khan):

are u there @baby?

OpenStudy (anonymous):

yea i am